OpenStudy (anonymous):

how to work out 4*(sqrt(1+[9/16]*x^2)) ?

OpenStudy (anonymous):

\[4 * \sqrt{1 + \frac{ 9 }{ 16 }x2} \]

OpenStudy (ivettef365):

do you have choices?

OpenStudy (anonymous):

i know the answer = \[\sqrt{(16+9x^2)}\] just no idea the steps to take to achieve it

OpenStudy (anonymous):

write \(4\) as \(\sqrt{16}\) so you can multiply inside the radical

OpenStudy (anonymous):

\[4 \times \sqrt{1 + \frac{ 9 }{ 16 }x2}=\sqrt{16}\times \sqrt{1 + \frac{ 9 }{ 16 }x2}\] \[=\sqrt{16(1+\frac{9}{16}x^2)}\] etc

OpenStudy (anonymous):

oh thank you very much! just when I think im on top of uni calculus, I get stumped on something that was probably taught in grade 2... thank you!

OpenStudy (anonymous):

yw i have to add i see no reason why one form is any nicer than the other, i.e. why you would want to turn \[4 \sqrt{1 + \frac{ 9 }{ 16 }x^2}\] into \[\sqrt{16+9x^2}\] or vice versa

OpenStudy (anonymous):

in my case it stems from finding the derivative of sinh^-1 [3x/4]

OpenStudy (anonymous):

and the multiple choice answer to the quiz was in the format of the latter and so I had to confirm that I could simplify to that form in someway before I could know I had the right answer :)